Calculate the MINIMUM value of the suitable range of fluid required for a 37 year old male patient weighing 240 pounds and 6 feet tall. This patient has suffered from severe trauma.
Answer #1

1. Initial resuscitation for severely injured patient based on the strategy of permissive hypovolemia, that is fluid resuscitation delivered until increasing B.P and make normotensive.

2. Crystaloids and colloids based transfusion used for severely injured patient.

Severe trauma fluid resuscitation calcuation:

Formula= 8 - 10ml/ kg/ hour.

Patient weight= 240pounds= 240/2.2= 108 kg.

So,

= 9ml/ 108/ hour

= 972ml per hour.